r - 5 square root of r - 36 = 0

solve equation over the set of complex numbers.

Answer by Nate(3500)

r - 5*sqrt(r - 36) = 0
r = 5*sqrt(r - 36)
r^2 = 25(r - 36)
r^2 = 25r - 900
r^2 - 25r + 900 = 0
r = (25 +- sqrt(625 - 3600))/2
r = (25 +- i*sqrt(2975))/2
r = (25 +- 5i*sqrt(119))/2
You may have to check those two answers, for one might be an exlusionary extraneous.